OpenStudy (anonymous):

Parabola Help. Write 3y^2+24y+72x-96=0 in standard form. I got (y+4)^2=-24(x-2) is that right? also what is the focus? line of symmetry? Directrix?

jimthompson5910 (jim_thompson5910):

you are correct 3y^2+24y+72x-96=0 is equivalent to (y+4)^2=-24(x-2)

jimthompson5910 (jim_thompson5910):

(y+4)^2=-24(x-2) (y-(-4))^2=-24(x-2) (y-(-4))^2=-4(6)(x-2) the equation is now in the form (y-k)^2 = -4p(x-h) where in this case h = 2 k =-4 p = 6 this means that 1) the parabola opens in a horizontal direction, either left or right because of the y^2 term 2) the parabola opens up to the left (since we have a negative -4 on the right side, if it was positive, then it would open up to the right) 3) the value of p determines how far the focus is from the vertex and it determines how far the directrix is from the vertex 4) the focus lies on the axis of symmetry

OpenStudy (anonymous):

got it thanks! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

tell me what you got for the focus and directrix

jimthompson5910 (jim_thompson5910):

oh and forgot to mention that (h,k) is the vertex, but I'm sure you know that already

OpenStudy (anonymous):

I got (-4,-4) and 8. vertex:(2,-4) of coarse.

jimthompson5910 (jim_thompson5910):

(-4,-4) is the focus, good x = 8 for the directrix (2,-4) is the vertex all 3 are correct

OpenStudy (anonymous):

yahoo! (:

jimthompson5910 (jim_thompson5910):

lol nice work